# Boltzmann Bridges

### v1

@@ VII.1 Model 1 Transition Rates
Let $\eta$ denote the number of maximum elements among the $x_{i}$. The probability to choose one of the maxima of $x_{i}$ is $\eta/d,$ and the probability of that choice increasing is $1/3$, so
@@
Thus $v$ is 00 unless there is only one maximum, and it equals $u$ otherwise. Thus, $u\geq v$. Now, we must compute the expectation value of $u,v$ over all sites on a given $r$, so as to compute $U(r),V(r),$ for $0<r<L$.
@@
We can prove via the binomial theorem that the above geometric structures defined with $w_{g}(r)$ partition the surface of constant $r$:
@@
where $V(r)$ is understood to be the averaged quantity, the probability of the radius decreasing during a timestep. $U(r)$ can be computed similarly: $U(r)=\frac{1}{w(r)}\times\sum_{\vec{x}}u(\vec{x}).$ Then,
@@
From here, it is simple to show that $U(r)\geq V(r).$ On the boundaries $r=0,r=L$, we have the probabilities
@@
independent of $d$. We have chosen the boundary condition that there is a maximum $L$. We make the simple choice that if a random walker were to increase in radius from $L$, it doesn’t move instead. Then, $V(L)$ is what is expected from Eq. (24), and $W(L)$ is such that $V(L)+W(L)=1$. This is in contrast to reweighting $v(L),w(L)$ based upon the convention that the random walker randomly selects uniformly over adjacent sites, which would result in $V(L)$ being greater than it is for the boundary conditions we have chosen. An example of the transition rates can be seen in Fig. 10.
@@ VII.2 Model 2 Conditional Distributions
We start by defining the following one-time conditioned probabilities:
@@
The fact that our dynamics is Markovian gives the following recursion relations among these quantities:
@@
Due only to Eq. (4), which are the symmetries of Eq. (26), the marginals $\mathbf{P}(s(t)=L)$ and $\mathbf{P}(s(t+k)=R)$ have to be time-independent, proven in section II.2. The only way for this to be possible while conserving probability is $\mathbf{P}(s(t)=L)=\mathbf{P}(s(t)=R)=1/2$, as expected.
We now seek to compute the general one-time conditioned distribution, Eq. (16). Denote the number of left to right transitions over $k$ timesteps to be $T^{LR}_{k}.$ First, we compute the probability that $T^{LR}_{k}=p$, given that $n_{L}(t)=n.$ Among the $n$ particles, choosing $p$ to switch gives a factor of $\binom{n}{p}$. The probability that any given of those particles is on the other side after the $k$ steps is $\alpha_{k}$. The probability that any of the rest of the $n-p$ particles are on their starting side is $\beta_{k}$. Multiplying the constituent probabilities and exponentiating to account for each particle, we have

### v2

@@ VII.1 Model 1 Transition Rates
Let $\eta$ denote the number of maximum elements among the $x_{i}$. The probability to choose one of them is $\eta/d,$ and the probability of that choice increasing is $1/3$, so
@@
Thus $v$ is 00 unless there is only one maximum, and it equals $u$ otherwise. Thus, $u\geq v$. Now, we must compute the expectation value of $u,v$ over all sites on a given $r$ in order to compute $U(r),V(r),$ for $0<r<L$.
@@
We can prove via the binomial theorem that these faces partition the surface of constant $r$:
@@
where $V(r)$ is the averaged quantity, the probability of the radius decreasing during a timestep. $U(r)$ is computed similarly: $U(r)=\frac{1}{w(r)}\times\sum_{\vec{x}}u(\vec{x}).$ Then,
@@
It is simple to show that $U(r)\geq V(r).$ On the boundaries $r=0,r=L$, we have the probabilities
@@
independent of $d$. We have chosen the boundary condition that there is a maximum $L$. We make the simple choice that if a random walker were to increase in radius from $L$, instead it doesn’t move. Then, $V(L)$ is what is expected from Eq. (24), and $W(L)$ changes so that $V(L)+W(L)=1$. This is in contrast to reweighting $v(L),w(L)$ based upon the convention that the random walker randomly selects uniformly over adjacent sites, which would result in $V(L)$ being greater. An example of the transition rates can be seen in Fig. 10.
@@ VII.2 Model 2 Conditional Distributions
We start with the one-time conditioned probabilities:
@@
Markovian dyanmics gives the following recursion relations among these quantities:
@@
Due only to Eq. (4), which are the symmetries of Eq. (26), the marginals $\mathbf{P}(s(t)=L)$ and $\mathbf{P}(s(t+k)=R)$ have to be time-independent, which we proved in Sec. II.2. The only way for this to be possible while conserving probability is $\mathbf{P}(s(t)=L)=\mathbf{P}(s(t)=R)=1/2$, as expected.
We now compute the general one-time conditioned distribution, Eq. (16). Denote the number of left to right transitions over $k$ timesteps to be $T^{LR}_{k}.$ First, we compute the probability that $T^{LR}_{k}=p$, given that $n_{L}(t)=n.$ Among the $n$ particles, choosing $p$ to switch gives a factor of $\binom{n}{p}$. The probability that any given of those particles is on the other side after the $k$ steps is $\alpha_{k}$. The probability that any of the rest of the $n-p$ particles are on their starting side is $\beta_{k}$. Multiplying the constituent probabilities and exponentiating to account for each particle, we have